Your small remodeling business has two work vehicles. One is a small passenger car used for job-site visits and for other general business purposes. The other is a heavy truck used to haul equipment. The car gets 25 miles per gallon (mpg). The truck gets 10 mpg. You want to improve gas mileage to save money, and you have enough money to upgrade one vehicle. The upgrade cost will be the same for both vehicles. An upgraded car will get 40 mpg; an upgraded truck will get 12.5 mpg. The cost of gasoline is $3.60 per gallon. Calculate the annual fuel savings, in gallons, for the truck and car assuming both vehicles are driven 11,000 miles per year. (Round your answers to 2 decimal places, e.g., 32.16.)

Truck per gallon_______?

Car per gallon_______?

Assuming an upgrade is a good idea in the first place, which one should you upgrade? Both vehicles are driven 11,000 miles per year. Truck or Car?
Answer #1

Number of miles per year = 11000

Current mileage of car = 25 mpg
=> Fuel required = 10000/25 = 400 gallons

New mileage of car = 40 mpg
=> Fuel Required = 10000/40 = 250 gallons

=> Fuel Savings in car = 400 - 250 = 150 gallons

Current mileage of truck = 10 mpg
=> Fuel required = 10000/10 = 1000 gallons

New mileage of truck = 12.5 mpg
=> Fuel Required = 10000/12.5 = 800 gallons

=> Fuel Savings in truck = 1000 - 800 = 200 gallons

Hence, we should upgrade the truck since the fuel savings in higher in case of a truck